### v1

# Coalescence in Markov chains

Abstract.
A Markov chain $X^{i}$ on a finite state space $S$ has transition matrix $P$ and initial state $i$. We may run the chains $(X^{i}:i\in S)$ in parallel, while insisting that any two such chains coalesce whenever they are simultaneously at the same state. There are $|S|$ trajectories which evolve separately, but not necessarily independently, prior to coalescence. What can be said about the number $k(\mu)$ of coalescence classes of the process, and what is the set $K(P)$ of such numbers $k(\mu)$, as the coupling $\mu$ of the chains ranges over couplings that are consistent with $P$? We continue earlier work of the authors (‘Non-coupling from the past’, in In and Out of Equilibrium 3, Springer, 2021) on these two fundamental questions, which have special importance for the ‘coupling from the past’ algorithm.
We concentrate partly on a family of couplings termed block measures, which may be viewed as couplings of lumpable chains with coalescing lumps. Constructions of such couplings are presented, and also of non-block measure with similar properties.
Key words and phrases:
Markov chain, grand coupling, coupling from the past, CFTP, lumpability, block measure, coalescence time, coalescence number, avoidance coupling
2010 Mathematics Subject Classification:
60J10, 60J22
1. Introduction
Finite-state space Markov chains form a key topic in probability theory, and they are taught in many undergraduate courses worldwide. They are considered to be well understood, and their theory fully established. Applications across science are significant and multifarious. In this paper we explore some interesting questions concerning coalescence that remain unresolved, and indeed in part unasked.
Throughout this paper, $S$ is a finite, non-empty set (without loss of generality, we may take $S=\{1,2,\dots,n\}$), and $P=(p_{i,j}:i,j\in S)$ is an irreducible stochastic matrix. Let $X=(X_{t}:t=0,1,\dots)$ be a Markov chain on the state space $S$ with transition matrix $P$. The long-term behaviour of $X$ is well known: if $X$ is (irreducible and) aperiodic, then $P$ has a unique invariant distribution ${\bm{\pi}}=(\pi_{j}:j\in S)$, and $\mathbb{P}(X_{t}=j)\to\pi_{j}$ for $j\in S$, as $t\to\infty$. For introductions to the theory of Markov chains, the reader is referred to the books [9, 10, 14].
For $i\in S$, denote by $X^{i}=(X_{t}^{i}:t=0,1,\dots)$ the Markov chain $X=(X_{t})$ conditioned on starting in state $i$, that is, conditioned on the event that $X_{0}=i$. In the current article we explore the simultaneous evolution of the chains $X^{i}$, $i\in S$, in the setting where any two paths stay together after they meet; that is, if $X^{i}_{t}=X^{j}_{t}$ then $X^{i}_{t+s}=X^{j}_{t+s}$ for all $s\in\mathbb{N}:=\{1,2,3,\dots\}$. For general $P$, multiple joint distributions of evolutions exist that are ‘consistent’ with $P$ in the sense that they have the required marginal distributions.
In this setting, we say that $X^{i}$ and $X^{j}$ coalesce (or just that $i$ and $j$ coalesce) if they meet (that is, if $X^{i}_{t}=X^{j}_{t}$ for some $t$). The principal issue investigated here is to determine the degree of coalescence of the family $(X^{i}:i\in S)$ as $t\to\infty$. This question is made more explicit in Section 3, and is connected to earlier work on so-called avoidance coupling; see Remark 3.4.
The current work is a development of earlier work of the authors, [8], directed mainly at coalescence in so-called ‘coupling from the past’ (CFTP). CFTP is an important technique for exact simulation from a given distribution on a state space $S$, and is related to the topic of ‘Monte Carlo Markov chains’ (MCMC). Whereas MCMC runs forwards in time, CFTP runs backwards. The CFTP algorithm is successful in situations where all the chains $(X^{i})$ coalesce. We explain this connection in Remark 3.2. Some background information from [8] is included in this work as an aid to the reader.
Here is a summary of the contents of this article. A ‘grand coupling’ is a coupling of the chains $(X^{i}:i\in S)$. Such couplings are defined in Section 2, where the special ‘independence coupling’ is introduced. In Theorem 2.3 we answer the question of when the independence coupling is the unique grand coupling. Section 3 is devoted to forward coalescence. Lumpable chains and block measures are introduced in Section 4, and the associated main results appear in Sections 5 and 6. A condition for a lumpable chain to give rise to a block measure is presented in Theorem 5.3; conversely, the existence of non-block measures is explored in Theorem 6.1 in the special case of the transition matrix with equally probable transitions. In the final Section 7, we initiate an investigation of the structure of the set of grand couplings corresponding to transition matrices.
We close with a remark. The general questions approached here seem quite fundamental to a complete theory of finite state-space Markov chains, and yet they have received only limited attention so far. This is reflected in the partial state of knowledge revealed in [8] and the current paper. This work may be viewed as a contribution to the theory of iterated random functions.
We write $\mathbb{N}$ for the set $\{1,2,\dots\}$ of natural numbers, and $\mathbb{P}$ for a generic probability measure.
2. Grand couplings of Markov chains
We shall consider only irreducible Markov chains on a finite state space $S$. Let $\mathcal{F}_{S}$ be the set of functions from $S$ to $S$, and let $\mathcal{P}_{S}$ be the set of irreducible stochastic matrices on $S$.
Definition 2.1 ([8]).
A probability measure $\mu$ on $\mathcal{F}_{S}$ is consistent with $P\in\mathcal{P}_{S}$, in which case we say that the pair $(P,\mu)$ is consistent, if
(2.1) $p_{i,j}=\mu\bigl(\{f\in\mathcal{F}_{S}:f(i)=j\}\bigr),\qquad i,j\in S.$
[…]
Let $P\in\mathcal{P}_{S}$. Any probability measure $\mu\in\mathcal{L}_{P}$ may be used as the basis of a coupling of the chains $(X^{i}:i\in S)$, as we now demonstrate. Let $F_{1},F_{2},\dots$ be random functions that are independent and distributed on $\mathcal{F}_{S}$ according to $\mu$, and set
(2.2) $X^{i}_{t}=F_{t}\circ F_{t-1}\circ\cdots\circ F_{1}(i),\qquad i\in S.$
(This is a convenient abuse of notation; more properly, the right side of (2.2) defines processes having the same law as the sequence $(X^{i})$.) For $t\in\{0,1,\dots\}$, let $Z_{t}=(X^{i}_{t}:i\in S)$. Then $Z=(Z_{t}:t=0,1,\dots)$ is a Markov chain with state space $S^{S}$, started in the state $(1,2,\dots,n)$. The ‘multichain’ $Z$ is in general reducible, since the number of distinct entries in $Z_{t+1}$ is less than or equal to that in $Z_{t}$.
By the consistency property (2.1), the sequence $X^{i}=(X_{t}^{i}:t=0,1,\dots)$ has the same law as $X$ conditioned on the event that $X_{0}=i$. Moreover, if $X_{t}^{i}=X_{t}^{j}$ then trivially $X_{t+s}^{i}=X_{t+s}^{j}$ for all $s\in\mathbb{N}$ (since each $F_{t}$ is a function); that is, once the paths from $i$ and $j$ meet, they stay together thenceforth.
A measure $\mu\in\mathcal{L}_{P}$ is sometimes called a grand coupling of $P$. It is elementary that $\mathcal{L}_{P}\neq\varnothing$, as indicated in the following example.
Example 2.2 (Independence coupling).
Let $P\in\mathcal{P}_{S}$, and let $\mu_{\text{\rm ind}}$ be given by the product form
$\mu_{\text{\rm ind}}(\{f\})=\prod_{i\in S}p_{i,f(i)},\qquad f\in\mathcal{F}_{S}.$
Then
$\mu_{\text{\rm ind}}(\{f:f(i)=j\})$ $=\sum_{f:f(i)=j}\prod_{u\in S}p_{u,f(u)}=p_{i,j}\sum_{f:f(i)=j}\prod_{u\neq i}p_{u,f(u)}$
$=p_{i,j}\sum_{(j_{1},\dots,j_{i-1},j_{i+1},\dots,j_{n})\in S^{n-1}}\prod_{u\neq i}p_{u,j_{u}}$
$=p_{i,j}\prod_{u\neq i}\biggl[\sum_{j_{u}\in S}p_{u,j_{u}}\biggr]=p_{i,j}\prod_{u\neq i}1,$
[…]
The next theorem contains a criterion for $\mu_{\text{\rm ind}}$ to be the unique grand coupling.
Theorem 2.3.
For $P\in\mathcal{P}_{S}$, we have $|\mathcal{L}_{P}|\geq 2$ if and only if $P$ has at least two rows each of which contains some entry lying in the open interval $(0,1)$.
Proof.
It can be useful to think in terms of the transition diagram of the chain, i.e., the labelled directed graph $G$ with vertex-set $S$, and with a directed edge from any $i$ to any $j$ such that $p_{i,j}>0$; such an edge is labelled with the value $p_{i,j}$. The in-degree (respectively, out-degree) of a state is the number of edges directed towards it (respectively, away from it).
Let $R$ be the set of rows that contain some entry in the interval $(0,1)$. Since $P$ is stochastic, such a row must contain at least two such entries.
(a) Let $|R|\leq 1$; we will show that the independence coupling is the unique grand coupling. We explain first the case $|R|=0$, for which every row contains a single non-zero entry $1$, and thus all edge-labels in $G$ are $1$. Each state $i$ has out-degree $1$, and we write $i,t_{i}\rangle$ for the unique directed edge leading from $i$. There are exactly $n=|S|$ edges in $G$. Since $P$ is irreducible, $G$ is a directed self-avoiding cycle of length $n$. Therefore, $\mathcal{L}_{P}$ contains only the probability measure $\mu=\delta_{f}$ that is supported on the single function $f$ defined by $f:i\mapsto t_{i}$ for $i\in S$. In this trivial situation, $\mu$ is the independence coupling $\mu_{\text{\rm ind}}$.
Suppose that $|R|=1$, and assume without loss of generality that the row labelled $1$ is the unique row containing an entry in $(0,1)$. Let $J=\{j:p_{1,j}\in(0,1)\}$. Edges of $G$ of the form $1,j\rangle$ for $j\in J$ have labels in $(0,1)$ and all other edges are labelled $1$. For $i\neq 1$, write $i,t_{i}\rangle$ for the unique edge directed away from $i$. We have
(2.3) $\mathbb{P}(X_{1}^{1}=j)$ $=p_{i,j}\quad$ $\text{for }j\in J,$
(2.4) $\mathbb{P}(X_{1}^{i}=t_{i})$ $=1\quad$ $\text{for }i\neq 1.$
There is a unique probability measure $\mu$ consistent with the above, namely
(2.5) $\mu(\{f\})=p_{1,f(1)}\prod_{i\neq 1}\delta_{f(i),t_{i}},\qquad f\in\mathcal{F}_{S},$
where $\delta$ is the Kronecker delta. This is simply the independence coupling. Therefore, $|\mathcal{L}_{P}|=1$.
(b) Conversely, suppose $|R|\geq 2$, and pick two rows lying in $R$, say the $i$th and $j$th with $i\neq j$. Find $r$, $s$ such that $p_{i,r}p_{j,s}>0$. There is no loss of generality for the proof that follows if we assume $i=1$ and $j=2$, and we assume this henceforth. We may assume further that
(2.6) $0<p_{2,s}\leq p_{1,r}<1.$
Before defining $\mu$ explicitly, we describe the stochastic evolution of the first time-step of the family $(X^{i}:i\in S)$ of Markov chains.
For $i\in\{2,3,\dots,n\}$, we let $X_{1}^{i}$ be chosen according to $P$, that is,
(2.7) $\mathbb{P}(X_{1}^{i}=j)=p_{i,j},\qquad j\in S.$
Furthermore, $Z:=\{X_{1}^{i}:i=2,3,\dots\}$ is a set of independent random variables.
We turn to $X_{1}^{1}$, which we take to be independent of $Z\setminus\{X_{1}^{2}\}$. Furthermore,
(2.8) $\mathbb{P}(X_{1}^{1}=r\mid X_{1}^{2}=s)$ $=1,$
(2.9) $\mathbb{P}(X_{1}^{1}=j\mid X_{1}^{2}\neq s)$ $=\begin{cases}\dfrac{p_{1,r}-p_{2,s}}{1-p_{2,s}}&\text{if }j=r,\\ \dfrac{p_{1,j}}{1-p_{2,s}}&\text{if }j\neq r.\end{cases}$
It is an exercise in conditional probability that the mass function of $X_{1}^{1}$ is $(p_{1,j}:j\in S)$. Equations (2.8)–(2.9) amount to a coupling of $X_{1}^{1}$ and $X_{1}^{2}$ under which (2.8) holds. Since $p_{1,r}<1$, (2.8) fails by (2.6) under the independence coupling $\mu_{\text{\rm ind}}$. Therefore, $\mu\neq\mu_{\text{\rm ind}}$ and hence $|\mathcal{L}_{P}|\geq 2$.
Equations (2.7)–(2.9) may be expressed by defining $\mu$ as follows:
(2.10) $\mu(\{f\})=\begin{cases}p_{2,s}\prod_{i\neq 1,2}p_{i,f(i)}&\text{ if }f(1)=r,\,f(2)=s,\\ \dfrac{p_{1,r}-p_{2,s}}{1-p_{2,s}}\prod_{i\neq 1}p_{i,f(i)}&\text{ if }f(1)=r,\,f(2)\neq s,\\ \dfrac{1}{1-p_{2,s}}\prod_{i}p_{i,f(i)}&\text{ if }f(1)\neq r,\,f(2)\neq s,\\ 0&\text{ otherwise}.\end{cases}$
One may check directly that $\mu\in\mathcal{L}_{P}$, but it is quicker to verify the probabilities implied by (2.7)–(2.9). ∎
Remark 2.4 (Random transition matrix).
By Theorem 2.3, a ‘typical’ transition matrix $P$ has multiple grand couplings. We make this statement more precise as follows. For given $S$, we model a ‘typical’ transition matrix as the $|S|\times|S|$ matrix with elements
$p_{i,j}=q_{i,j}/Q_{i},$
[…]
We will make frequent use of the following well known fact due to Doeblin [5] (see also [9, p. 260]).
Theorem 2.5 ([5]).
Let $(X^{i}:i\in S)$ be independent Markov chains on a finite state space $S$ with common irreducible, aperiodic transition matrix $Q$, with $X^{i}_{0}=i$ a.s. for each $i\in S$. For $i,j\in S$, there exists a.s. a finite time $T$ such that $X^{i}_{T}=X^{j}_{T}$.
The following classical result of Birkhoff and von Neumann will be useful later.
Theorem 2.6 ([3, 13]).
A stochastic matrix $P$ on the finite state space $S$ is doubly stochastic if and only if it lies in the convex hull of the set $\Pi_{S}$ of permutation matrices.
3. Coalescence of trajectories
Consider a Markov chain $X$ on the state space $S=\{1,2,\dots,n\}$ with irreducible transition matrix $P$, and let $\mu\in\mathcal{L}_{P}$ be a grand coupling. We consider the coalescence of the chains $X^{i}$ in this section, and begin with some notation from [8].
As in Section 2 (see (2.2)), let $F_{1},F_{2},\dots$ be independent random functions distributed according to $\mu$, and set $X^{i}_{t}={\overset{{}_{\shortrightarrow}}{F}}_{t}(i)$ where $i\in S$ and
(3.1) ${\overset{{}_{\shortrightarrow}}{F}}_{t}:=F_{t}\circ F_{t-1}\circ\cdots\circ F_{1}.$
Then $(X^{i}:i\in S)$ is a family of coupled Markov chains running forwards in time, each having transition matrix $P$, and such that $X^{i}$ starts in state $i$.
For $i,j\in S$, let
(3.2) $T_{i,j}=\inf\{t:X^{i}_{t}=X^{j}_{t}\}.$
We say that $i$ and $j$ coalesce (and write $i\sim j$) if $T_{i,j}<\infty$. The forward coalescence time is given by
(3.3) $T$ $=\inf\bigl\{t:{\overset{{}_{\shortrightarrow}}{F}}_{t}(\cdot)\text{ is a constant function}\bigr\}=\sup_{i,j\in S}T_{i,j}.$
We say that coalescence occurs if $\mathbb{P}(T<\infty)=1$.
Lemma 3.1.
The relation $\sim$ is a (random) equivalence relation.
Proof.
It suffices to prove the transitivity of $\sim$. Let $i\sim j$ and $j\sim k$. Since $S$ is finite, there exists $T$ such that $X_{t}^{i}=X_{t}^{j}=X_{t}^{k}$ for $t\geq T$. The claim follows ∎
The equivalence classes of $\sim$ are termed the coalescence classes of $(X^{i}:i\in S)$. The number of coalescence classes is a.s. constant (see [8, Lemma 2]), and we denote this number by $k(\mu)$ and call it the coalescence number of $\mu$. We define
(3.4) $K(P)=\{k(\mu):\mu\in\mathcal{L}_{P}\}.$
Remark 3.2 (Coupling from the past (CFTP)).
CFTP is a prominent algorithm introduced by Propp and Wilson [15, 16, 18] for perfect simulation from the invariant distribution of an irreducible, aperiodic Markov chain on a finite state space. The CFTP algorithm may be defined by replacing the function ${\overset{{}_{\shortrightarrow}}{F}}_{t}$ in (3.1) by the function ${\overset{{}_{\shortleftarrow}}{F}}_{t}:=F_{1}\circ F_{2}\circ\dots\circ F_{t}$. The backward coalescence time is defined by
(3.5) $C=\inf\bigl\{t:{\overset{{}_{\shortleftarrow}}{F}}_{t}(\cdot)\text{ is a constant function}\bigr\},$
and backward coalescence is said to occur if $\mathbb{P}(C<\infty)$. On the event $\{C<\infty\}$, ${\overset{{}_{\shortleftarrow}}{F}}_{C}$ may be regarded as a random state, and the main theorem of CFTP asserts that, if backward coalescence occurs, then ${\overset{{}_{\shortleftarrow}}{F}}_{C}$ is distributed as the invariant distribution of the transition matrix $P$.
[…]
The following two questions are fundamental to understanding coalescence.
Question 3.3.
  • 1.
    Can we determine the set $K(P)$ for given $P$?
  • 2.
    Which $\mu\in\mathcal{L}_{P}$ have $k(\mu)=1$?
We shall see in Theorem 3.13 that the independence coupling $\mu_{\text{\rm ind}}$ of Example 2.2 satisfies $k(\mu_{\text{\rm ind}})\leq k(\mu)$ for all $\mu\in\mathcal{L}_{P}$, and moreover $k(\mu_{\text{\rm ind}})=1$ if and only if $P$ is aperiodic.
Henceforth, expressions involving the word ‘coalescence’ shall refer to forward coalescence. Let $\mu$ be a probability measure on $\mathcal{F}_{S}$, and let $\text{\rm supp}(\mu)$ denote the support of $\mu$. Let $F=(F_{s}:s\in\mathbb{N})$ be a vector of independent and identically distributed random functions, each with law $\mu$. The law of $F$ is the product measure ${\bm{\mu}}=\prod_{i\in\mathbb{N}}\mu$.
Remark 3.4 (Avoidance coupling).
Let $P\in\mathcal{P}_{S}$ and let
$k_{\text{\rm max}}=k_{\text{\rm max}}(P):=\max\{k(\mu):\mu\in\mathcal{L}_{P}\}.$
[…]
Recall that, since $P$ is finite and irreducible, it has a unique invariant distribution (see, for example, [9, Thm 6.4.3]).
Theorem 3.5.
Let $P\in\mathcal{P}_{S}$ have (unique) invariant distribution $\bm{\pi}$}.
  • (a)
    Let $m\in\mathbb{N}$, and suppose there exists $s\in S$ with $\pi_{s}>1/m$. Then $k_{\text{\rm max}}<m$.
  • (b)
    Let $m\in\mathbb{N}$, and suppose there exists $s\in S$ such that $p_{i,s}>1/m$ for all $i\in S$, then $k_{\text{\rm max}}<m$.
Proof.
(a) Assume $\pi_{s}>1/m$, and let $i_{1},i_{2},\dots,i_{m}$ be distinct elements of $S$. For each $k\in\{1,2,\dots,m\}$, there is asymptotic density $\pi_{s}$ of times $n$ at which $X_{n}^{i_{k}}=s$. Since $\pi_{s}>1/m$, there exist (a.s.) distinct $i,j\in\{i_{1},i_{2},\dots,i_{m}\}$ such that there is a strictly positive density of times $n$ at which $X_{n}^{i}=X_{n}^{j}=s$. Therefore, $k(\mu)<m$ for all $\mu\in\mathcal{L}_{P}$.
(b) Assume the given condition holds. Then
(3.6) $\pi_{s}=\sum_{i\in S}\pi_{i}p_{i,s}>\frac{1}{m}.$
The conclusion holds by part (a). ∎
Remark 3.6.
The condition of part (b) may be changed slightly, as follows. Suppose $p_{i,s}>1/m$ for all $i\in S$ with $i\neq s$. Then (3.6) becomes
$\pi_{s}\geq\sum_{i\neq s}\pi_{i}p_{i,s}>\frac{1}{m}(1-\pi_{s}).$
[…]
Remark 3.7 (Random transition matrix).
[…]
Question 3.8.
Is it true that $K(P)=\{1\}$ for $\mathbb{Q}$-a.e. $P$?
Whereas $k(\mu)$ is a.s. constant, the coalescence classes of $\sim$ need not themselves be a.s. constant. Here is an example of this, preceded by some notation.
Definition 3.9.
[…]
Example 3.10.
Take $S=\{1,2,3,4\}$ and any consistent pair $(P,\mu)$ with
$\text{\rm supp}(\mu)=\{f_{1,1},f_{1,2},f_{2,1},f_{2,2}\}$
where
$f_{1,1}=(1212),\quad f_{1,2}=(1221),\quad f_{2,1}=(3434),\quad f_{2,2}=(3443).$
[…]
$12341234$
Figure 3.1. Diagrammatic representations of the four functions $f_{i,j}$ of Example 3.10.
Question 3.11.
For given $P\in\mathcal{P}_{S}$, is the set of cardinalities (possibly with repetition) of coalescing classes a deterministic set?
A probability measure $\mu$ on $\mathcal{F}_{S}$ may be written in the form
(3.7) $\mu=\sum_{f\in\mathcal{F}_{S}}\alpha_{f}\delta_{f},$
where $\bm{\alpha}$} is a probability mass function on $\mathcal{F}_{S}$ with support $\text{\rm supp}(\mu)$, and $\delta_{f}$ is the Dirac delta-mass on the point $f\in\mathcal{F}_{S}$. Thus, $\alpha_{f}>0$ if and only if $f\in\text{\rm supp}(\mu)$. We denote by $M(\mathcal{F}_{S})$ the set of all probability measures $\mu$ of the form (3.7).
In advance of stating an extension of [8, Lemma 3], we remind the reader of the definition of a cyclic class.
Definition 3.12.
Consider an irreducible Markov chain on the state space $S$ with transition matrix $P$ and period $d$. There exists a unique partition of $S$ into $d$ classes $S_{0},S_{1},\dots,S_{d-1}$ such that
$\sum_{j\in S_{r+1}}p_{i,j}=1,\qquad i\in S_{r},\ r=0,1,\dots,d-1,$
[…]
See [4, Sect. 2.3.2] for further details of cyclic classes.
Theorem 3.13.
  • (a)
    We have that $1\in K(P)$ if and only if $P\in\mathcal{P}_{S}$ is aperiodic. In this case, $k(\mu_{\text{\rm ind}})=1$.
  • (b)
    If $P\in\mathcal{P}_{S}$ has period $d$, then the coalescence classes of $\mu_{\text{\rm ind}}$ are a.s. the cyclic classes of $P$ and therefore $k(\mu_{\text{\rm ind}})=d$. Hence, $d\in K(P)$, and moreover $d\leq k$ for all $k\in K(P)$.
Consider the particular case when all entries of $P$ are strictly positive. Since $P$ is aperiodic, we have by Theorem 3.13(a) that $1\in K(P)$. However, $K(P)$ may be larger than a singleton; see the forthcoming Theorem 6.1.
Proof of Theorem 3.13.
(a) This is found at the proof of [8, Lemma 3], but is included here for clarity. The independence coupling of Example 2.2 gives rise to $n=|S|$ chains with transition matrix $P$, starting from $1,2,\dots,n$, respectively, that evolve independently until they meet. If $P$ is aperiodic (and, by assumption, irreducible) then all $n$ chains meet a.s. in finite time. The last holds by Theorem 2.5, since any paths that meet will remain together thenceforth under this coupling.
Conversely, if $P$ is periodic and $p_{i,j}>0$ then $i\neq j$, and $i$ and $j$ can never coalesce, implying $1\notin K(P)$.
(b) Suppose $P$ has period $d\geq 2$, and let $S_{0},S_{1},\dots,S_{d-1}$ be the cyclic classes of $P$ (see Definition 3.12). For $r\neq s$, states $i\in S_{r}$ and $j\in S_{s}$ do not coalesce for any $\mu\in\mathcal{L}_{P}$. Now, $P^{d}$ has a block diagonal form with (transition) matrices $E_{0},E_{1},\dots,E_{d-1}$ along its diagonal. Each $E_{r}$ is irreducible and aperiodic. By the argument above, subject to the independence coupling $\mu_{\text{\rm ind}}$, any two states in any given $S_{r}$ coalesce a.s.
For clarity, we give some more details of the last step. Fix $r$ and let $i,j\in S_{r}$. Let $Z^{k}=(Z_{t}^{k}:t=0,1,\dots)$ be a copy of $X^{k}$, and suppose the chains $(Z^{k}:k\in S_{r})$ evolve independently (unlike the sequence $(X^{k})$ whose members coalesce when they meet). Now, $Z^{k,d}:=(Z^{k}_{md}:m=0,1,\dots)$ has irreducible, aperiodic transition matrix $E_{r}$. By Theorem 2.5 there exists a.s. some time $T<\infty$ at which $Z^{i}_{Td}=Z^{j}_{Td}$. Therefore, there is some earliest time $U\leq Td$ at which $Z^{i}_{U}=Z^{j}_{U}$. From the pair $Z^{i}$, $Z^{j}$ we construct $X^{i}$, $X^{j}$ by
$(X^{i}_{t},X^{j}_{t})=\begin{cases}(Z^{i}_{t},Z^{j}_{t})&\text{for }t\leq U,\\ (Z^{i}_{t},Z^{i}_{t})&\text{for }t>U.\end{cases}$
Since the pair $(X^{i},X^{j})$ has the same law as under $\bm{\mu}$}, this proves that $i$ and $j$ coalesce a.s. under $\bm{\mu}$}. This holds for all pairs of distinct states in $S_{r}$, and two states stick together after they coaleasce. The second claim of the theorem is proved.
The minimality of $d$ holds since no two states in distinct cyclic classes may coalesce. ∎
4. Lumpability and block measures
The notion of lumpability was introduced in 1963 by Kemeny and Snell, [11]. A Markov chain $X$ with finite state space $S$ is called lumpable if there exists a partition $\mathcal{S}$ of $S$ such that the projection of $X$ onto $\mathcal{S}$ is itself a Markov chain.
Here is a more precise definition. A partition $\mathcal{S}=\{S_{1},S_{2},\dots,S_{\ell}\}$ of $S$ is called trivial if $|\mathcal{S}|\in\{1,|S|\}$, (i.e., if either $\mathcal{S}=\{S\}$ or $\mathcal{S}=\{\{s\}:s\in S\}$). Elements of the partition $\mathcal{S}$ are called blocks. The chain $X$ is called $\mathcal{S}$-lumpable if the sequence $(W_{t}:t=0,1,\dots)$, defined by
(4.1) $W_{t}=j\quad\text{if}\quad X_{t}\in S_{j},$
forms a Markov chain. The process $X$ is called lumpable if it is $\mathcal{S}$-lumpable for some non-trivial partition $\mathcal{S}$.
There is a limited literature on lumpable chains, for which the reader may consult, for example, [6, 7, 12] and [9, Exer. 6.1.13]. Kemeny and Snell [11] proved a necessary and sufficient condition for $X$ to be $\mathcal{S}$-lumpable, namely the following.
Theorem 4.1 ([11]).
Let $P=(p_{i,j}:i,j\in S)$. Let $\mathcal{S}$ be a partition of $S$, and let
(4.2) $\lambda_{r,s}^{(i)}=\sum_{j\in S_{s}}p_{i,j},\quad i\in S_{r}.$
A Markov chain $X$ with transition matrix $P$ is $\mathcal{S}$-lumpable if and only if,
(4.3) for every $r$, $s$, we have that $\lambda_{r,s}:=\lambda_{r,s}^{(i)}$ is constant for $i\in S_{r}$.
A stochastic matrix $P\in\mathcal{P}_{S}$ is called $\mathcal{S}$-lumpable if (4.3) holds (where the $\lambda_{r,s}^{(i)}$ are given in (4.2)). For such a pair $P$ and $\mathcal{S}$, let $\Lambda$ be the $\ell\times\ell$ matrix $(\lambda_{r,s}:1\leq r,s\leq\ell)$. This $\Lambda$ is the transition matrix of the ‘block process’ $W$ of (4.1).
The evolution of an $\mathcal{S}$-lumpable chain $X$ may be given in two stages. Suppose $X_{t}=i\in S_{r}$; then $X_{t+1}$ is given as follows.
  • (a)
    Select a random block $B$ with mass function $\mathbb{P}(B=S_{s})=\lambda_{r,s}$.
  • (b)
    Conditional on $B=S_{s}$, choose $X_{t+1}$ with mass function
    $\mathbb{P}(X_{t+1}=j\mid B=S_{s})=p_{i,j}/\lambda_{r,s}.$
We address a certain sub-category of lumpable chains here, namely those for which $\Lambda$ is doubly stochastic. By Theorem 2.6, such $\Lambda$ may be expressed as a convex combination of permutation matrices.
Recall the set $M(\mathcal{F}_{S})$ of probability measures of the form (3.7).
Definition 4.2 ([8]).
Let $\mu\in M(\mathcal{F}_{S})$. For a partition $\mathcal{S}=\{S_{r}:r\in I\}$ of $S$ with index set $I=\{1,2,\dots,\ell\}$, we call $\mu$ an $\mathcal{S}$-block measure if
  • (a)
    for $f\in\text{\rm supp}(\mu)$, there exists a unique permutation $\pi=\pi_{f}$ of $I$ such that, for $r\in I$, $fS_{r}\subseteq S_{\pi(r)}$, and
  • (b)
    $k(\mu)=\ell$.
[…]
Since any two states in distinct blocks cannot coalesce under a block measure, the condition $k(\mu)=\ell$ implies that
(4.4) for $r\in I$ and $i,j\in S_{r}$, the pair $i$, $j$ coalesce a.s.,
so that the coalescence classes of the coalescence relation $\sim$ are a.s. the blocks $S_{1},S_{2},\dots,S_{\ell}$.
By combining the definitions of lumpability and block measures we arrive at a necessary condition for $\mu$ to be an $\mathcal{S}$-block measure. The proof may be found in [8].
Theorem 4.3 ([8, Thm 6]).
Let $S$ be a non-empty, finite set, let $P\in\mathcal{P}_{S}$, and let $\mathcal{S}=\{S_{r}:r\in I\}$ be a partition of $S$ with index set $I=\{1,2,\dots,\ell\}$. For $i\in S_{r}$, let
(4.5) $\lambda_{r,s}^{(i)}=\sum_{j\in S_{s}}p_{i,j}.$
  • (a)
    If $\mu\in\mathcal{L}_{P}$ is an $\mathcal{S}$-block measure, then, for $r,s\in I$,
    (4.6) $\text{$\lambda_{r,s}:=\lambda_{r,s}^{(i)}$ is constant for $i\in S_{r}$}.$
  • (b)
    The ‘block-transition matrix’ $\Lambda=(\lambda_{r,s}:1\leq r,s\leq\ell)$ is irreducible. Moreover, $\Lambda$ is doubly stochastic, which may be expressed as
    (4.7) $\sum_{i\in S}\sum_{j\in S_{s}}\frac{1}{|S_{r(i)}|}p_{i,j}=1,\qquad r,s\in I,$
    where $r(i)$ is the index $r\in I$ such that $i\in S_{r}$.
Whether or not $\mu\in M(\mathcal{F}_{S})$ is a block measure turns out to depend on whether or not its coalescing classes are deterministic, that is, a.s. constant. (Recall the definition of coalescing class after Lemma 3.1.) We state this as a theorem.
Theorem 4.4.
A probability measure $\mu\in M(\mathcal{F}_{S})$ is a block measure if and only if its coalescing classes $\mathcal{C}=\{C_{1},C_{2},\dots,C_{\ell}\}$ are a.s. constant. If the last holds, $\mu$ is a $\mathcal{C}$-block measure.
Proof.
If $\mu\in M(\mathcal{F}_{S})$ is an $\mathcal{S}$-block measure then (recall (4.4) and the preceding discussion) the coalescing classes are a.s. the elements of $\mathcal{S}$.
Conversely, let the coalescing classes be $C_{1},\dots,C_{\ell}$, and assume they are a.s. constant. We claim that Definition 4.2 is satisfied with $\mathcal{S}=\mathcal{C}$. Clearly $k(\mu)=\ell$, so (b) of Definition 4.2 holds. Let $f\in\text{\rm supp}(\mu)$, and let $i_{1},i_{2}\in C_{j}$. Since $i_{1}$ and $i_{2}$ a.s. coalesce, $f(i_{1})$ and $f(i_{2})$ a.s. coalesce, so $f(i_{1})\in C_{s}\iff f(i_{2})\in C_{s}$. Suppose instead that $i_{1}\in C_{j_{1}}$ and $i_{2}\in C_{j_{2}}$ where $j_{1}\neq j_{2}$. This implies that $i_{1}$ and $i_{2}$ cannot coalesce, and hence $f(i_{1})\in C_{s}$ implies $f(i_{2})\notin C_{s}$.
We have shown that $f$ permutes classes, which verifies Definition 4.2(a) with $\mathcal{S}=\mathcal{C}$, and hence completes the proof. ∎
Example 4.5.
Here is an illustration of Theorem 4.4. Consider the transition matrix
$P=\begin{pmatrix}\frac{1}{2}&0&\frac{1}{2}&0\\ 0&\frac{1}{2}&0&\frac{1}{2}\\ 0&\frac{1}{2}&\frac{1}{2}&0\\ \frac{1}{2}&0&0&\frac{1}{2}\end{pmatrix}$,
and note that it is irreducible and aperiodic.
  • (a)
    Let $\mu_{f}$ put mass $\frac{1}{2}$ on each of the two permutations
    $f_{1}=(1234),\qquad f_{2}=(3421).$
    (Recall Definition 3.9.) It is easily checked that $\mu_{f}\in\mathcal{L}_{P}$ is a (trivial) block measure with blocks $\{1\}$, $\{2\}$, $\{3\}$, $\{4\}$, so that $k(\mu_{f})=4$.
  • (b)
    Let $\mu_{g}$ put mass $\frac{1}{2}$ on each of the two functions
    $g_{1}=(3434),\qquad g_{2}=(1221).$
[…]
5. Existence of block measures
This section is concerned with the existence of block measures. The main question of interest is, for what state space $S$, partition $\mathcal{S}$, and matrix $P\in\mathcal{P}_{S}$ does there exist an $\mathcal{S}$-block measure $\mu\in\mathcal{L}_{P}$? Let $\mathcal{C}=\mathcal{C}(\mu)$ denote the set of coalescing classes of $\mu$. In general $\mathcal{C}(\mu)$ is random, and the support of $\mathcal{C}(\mu)$ depends only on $\text{\rm supp}(\mu)$. By Theorem 4.4, finding an $\mathcal{S}$-block measure $\mu\in\mathcal{L}_{P}$ is equivalent to finding $\mu\in\mathcal{L}_{P}$ for which $\mathcal{C}=\mathcal{S}$ a.s.
By Theorem 3.13, for every $P\in\mathcal{P}_{S}$ there exists a partition $\mathcal{S}$ of $S$ (comprising the cyclic classes of $P$) such that the independence coupling $\mu_{\text{\rm ind}}$ is an $\mathcal{S}$-block measure. Recall the probability measure $\mathbb{Q}$ in Remark 2.4.
Lemma 5.1.
For $\mathbb{Q}$-a.e. $P\in\mathcal{P}_{S}$, every block measure is trivial.
Proof.
We need only consider cases with $|S|\geq 3$. If there exists a non-trivial block measure, then there exists $T\subseteq S$ such that $2\leq|T|\leq|S|-1$ and, by (4.5)–(4.6),
$\sum_{j\in T}p_{i,j}\text{ is constant for }i\in T.$
Let $Z_{i}:=\sum_{j\in T}p_{i,j}$. Then $Z_{1},Z_{2},\dots$ are independent with a common absolutely continuous distribution. Therefore, $\mathbb{Q}(Z_{i}=Z_{k})=0$ for all distinct pairs $i$, $k$, whence the set of such $P$ has $\mathbb{Q}$-measure 0. ∎
By Lemma 5.1, for every $S$, for $\mathbb{Q}$-a.s. $P\in\mathcal{P}_{S}$, and for every non-trivial $\mathcal{S}$, there exists no $\mathcal{S}$-block measure. In contrast, the following holds.
Theorem 5.2.
For $S=\{1,2,\dots,n\}$ and any partition $\mathcal{S}$ of $S$ there exists $\mu\in M(\mathcal{F}_{S})$ such that $\mu$ is an $\mathcal{S}$-block measure.
Proof.
Fix $S$ and $\mathcal{S}=\{S_{r}:r\in I\}$. Consider the set $\widehat{\mathcal{F}}$ of functions $f$ such that
  • (i)
    for all $r$, $f(i)=f(j)$ for every $i,j\in S_{r}$, and
  • (ii)
    there exists a permutation $\pi=\pi_{f}$ of $I$ such that, for $r\in I$, $fS_{r}\subseteq S_{\pi(r)}$.
For example, if $\mathcal{S}=\{\{1,2\},\{3,4,5\}\}$ then $\widehat{\mathcal{F}}$ consists of the 12 functions denoted
$(11333),(11444),(11555),(22333),(22444),(22555),$
$(33111),(44111),(55111),(33222),(44222),(55222),$
in the notation of Definition 3.9.
Let $\mu$ be a probability measure with $\text{\rm supp}(\mu)=\widehat{\mathcal{F}}$. Then $p_{i,j}:=\sum_{f:f(i)=j}\mu(\{f\})>0$ for every pair $i$, $j$, whence $P=(p_{i,j})$ is irreducible. The claim follows by assumptions (i) and (ii), and Definition 4.2. ∎
Henceforth, for certain $P$, $\mathcal{S}$, we will present a natural measure $\mu\in\mathcal{L}_{P}$, and then determine conditions under which it is an $\mathcal{S}$-block measure.
Let the pair $P\in\mathcal{P}_{S}$ and $\mathcal{S}=\{S_{r}:r\in I\}$ (a partition of the state space $S$ with $I=\{1,2,\dots,\ell\}$) satisfy (4.6) and be such that
(5.1) $\Lambda$ is doubly stochastic.
By Theorem 4.3 these two conditions are necessary for the existence of an $\mathcal{S}$-block measure $\mu\in\mathcal{L}_{P}$.
Since $P$ is assumed irreducible, the stochastic matrix $\Lambda$ is irreducible also. By (5.1) and Theorem 2.6, we may find a measure $\rho\in\mathcal{L}_{\Lambda}$ supported on a subset of the set of permutations of $I$, and we let $\Pi$ be a random permutation with law $\rho$. (Note that $\rho$ is not generally unique.) Let $i\in S_{r}$ and $j\in S_{s}$. In order for our forthcoming $\mu$ to be consistent with $P$ (recall Definition 2.1) we need that
(5.2) $\sum_{f:i\mapsto j}\mu(\{f\})=p_{i,j}.$
In order for $i$ to be mapped to $j$, it is necessary that $\Pi(r)=s$; the last occurs with probability $\lambda_{r,s}$. Conditional on $\Pi$, we shall then map the states independently in such a way as to obtain (5.2).
More precisely, conditional on $\Pi$, let $Z=(Z_{i}:i\in S)$ be independent random variables such that
(5.3) $\mathbb{P}(Z_{i}=j\mid\Pi)=\begin{cases}p_{i,j}/\lambda_{r,s}&\text{if }S_{r}\ni i,\ S_{s}\ni j,\ \Pi(r)=s,\\ 0&\text{otherwise},\end{cases}$
and let $\mu$ be the law of $Z$. Thus, $\mu\in M(\mathcal{F}_{S})$ is given by
(5.4) $\mu(\{f\})=\mathbb{E}\left[\prod_{i\in S}\mathbb{P}(Z_{i}=f(i)\mid\Pi)\right],$
where the expectation is over the random permutation $\Pi$. We call $\mu$ the $(P,\mathcal{S},\rho)$-product measure.
Next we check that $\mu\in\mathcal{L}_{P}$. Let $i\in S_{r}$ and $j\in S_{s}$. By (5.3)–(5.4),
$\sum_{f:i\mapsto j}\mu(\{f\})=\mathbb{P}(Z_{i}=j)=\lambda_{r,s}\cdot\dfrac{p_{i,j}}{\lambda_{r,s}}=p_{i,j},$
as required. By the definition of $\mu$, no two states in different sets of the partition $\mathcal{S}$ may coalesce. Thus, to determine whether the $(P,\mathcal{S},\rho)$-product measure $\mu\in\mathcal{L}_{P}$ is an $\mathcal{S}$-block measure it remains to check whether $k(\mu)=\ell=|\mathcal{S}|$, which is to say that, for all $r$, all states in $S_{r}$ coalesce (recall (4.4)).
Recall that $\mathcal{C}$ denotes the (possibly random) set of coalescing classes, and that $i\sim j$ if there exists $C\in\mathcal{C}$ such that $i,j\in C$.
Theorem 5.3.
Let $P\in\mathcal{P}_{S}$, and let $\mathcal{S}=\{S_{r}:r\in I\}$ be a partition of $S$ with index set $I=\{1,2,\dots,\ell\}$. Assume (4.6) and (5.1) hold. The $(P,\mathcal{S},\rho)$-product measure $\mu$ satisfies $k(\mu)=\ell$ if and only if [^2] One may work with any given value of $r$ in (5.5), and we have chosen $r=1$ for concreteness. Condition (5.5) is similar to the sufficient condition of [17, Thm 6] for the quenched ergodicity of a Markov chain with random transition matrices. Theorem 5.3 corrects an error in [8, Thm 6].
(5.5) $\mathbb{P}(i\sim j)>0,\qquad i,j\in S_{1}.$
Proof.
Suppose the conditions of the theorem hold, and also (5.5). Let $r\in I$. We will show that
(5.6) $\mathbb{P}(i\sim j)=1,\qquad i,j\in S_{r}.$
The claim $k(\mu)=\ell$ follows since there are finitely many pairs $i,j\in S_{r}$ and indices $r$.
Recall from (3.2) that $T_{i,j}=\inf\{t\geq 0:X_{t}^{i}=X_{t}^{j}\}$, so that $i\sim j$ if and only if $T_{i,j}<\infty$.
We first prove (5.6) with $r=1$. Since $|S_{1}|<\infty$, by (5.5) there exists $\epsilon>0$, and $M<\infty$ such that
(5.7) $\mathbb{P}(T_{i,j}\leq M)>\epsilon,\qquad i,j\in S_{1}.$
Let $i,j\in S_{1}$. Either $T_{i,j}\leq M$ or not. Assume that $T_{i,j}>M$. We continue the two chains from time $M$ with initial states $X_{M}^{i}$ and $X_{M}^{j}$ until the next epoch ($M+K$, say) at which these two processes lie in $S_{1}$; note that $\mathbb{P}(K<\infty)=1$ since $\Lambda$ is irreducible. Having arrived back in $S_{1}$, we apply the argument above to deduce that coalescence occurs by time $2M+K$ with (conditional) probability at least $\epsilon$.
It follows that
$\mathbb{P}(T_{i,j}>2M+K)$ $=\mathbb{P}(T_{i,j}>2M+K\mid T_{i,j}>M)\mathbb{P}(T_{i,j}>M)$
$\leq(1-\epsilon)^{2}.$
By iteration of this argument, $\mathbb{P}(T_{i,j}<\infty)=1$, and (5.6) follows with $r=1$.
For $r\neq 1$ and $i,j\in S_{r}$, since $\Lambda$ is irreducible, a.s. $S_{r}$ is eventually mapped to $S_{1}$. At this point, $i$ and $j$ have both been mapped to elements of $S_{1}$, so (if they have not already done so) they will almost surely coalesce by (5.6) with $r=1$. This verifies (5.6) for general $r$ as required.
Conversely, if there exist $i,j\in S_{1}$ such that $\mathbb{P}(i\sim j)=0$ then it is necessarily the case that $k(\mu)\geq\ell+1$. ∎
Example 5.4.
Here is an illustration of Theorem 5.3. As there (with $\ell=2$ for simplicity) we let $P\in\mathcal{P}_{S}$, $\mathcal{S}=\{S_{1},S_{2}\}$, and we assume that (4.6) and (5.1) hold.
For $r=1,2$ we write $S_{r}=\{x_{r,j}:1\leq j\leq m_{r}\}$. Ordering the elements of $S$ as $(x_{1,1}.\dots,x_{1,m_{1}},x_{2,1},\dots,x_{2,m_{2}})$, we may express the $(m_{1}+m_{2})\times(m_{1}+m_{2})$ matrix $P$ in the form
$P=\begin{pmatrix}A&B\\ C&D\end{pmatrix}$,
where $A$ is an $m_{1}\times m_{1}$ matrix and $D$ is an $m_{2}\times m_{2}$ matrix.
The $2\times 2$ matrix $\Lambda$ is doubly stochastic and irreducible, and may (by Theorem 2.6) be expressed in the form
(5.8) $\Lambda=\begin{pmatrix}\lambda_{1,1}&\lambda_{1,2}\\ \lambda_{2,1}&\lambda_{2,2}\end{pmatrix}=\alpha I+(1-\alpha)\overline{I}$
where $\alpha=\lambda_{1,1}=\lambda_{2,2}$ and
$I=\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\qquad\overline{I}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$.
In other words, the measure $\rho$ (which is unique since $\Lambda$ is a $2\times 2$ matrix) puts mass $\alpha$ on the identity permutation, and mass $1-\alpha$ on the ‘interchange’ permutation. We let $(\Pi_{i}:i\in\mathbb{N})$ be independent permutations with common law $\rho$, and let $\mu$ be the $(P,\mathcal{S},\rho)$-product measure.
We have shown that
$\Lambda=\begin{pmatrix}\alpha&1-\alpha\\ 1-\alpha&\alpha\end{pmatrix}$.
Thus, the row sums of $A$ and $D$ (respectively, $B$ and $C$) are all $\alpha$ (respectively, $1-\alpha$) by (4.5) and (4.6)). Therefore, $P$ is a mixture of two stochastic matrices $P_{1}$ and $P_{2}$,
$P=\alpha P_{1}+(1-\alpha)P_{2},$
where
$P_{1}=\frac{1}{\alpha}\begin{pmatrix}A&0\\ 0&D\end{pmatrix},\qquad P_{2}=\frac{1}{1-\alpha}\begin{pmatrix}0&B\\ C&0\end{pmatrix}$.
Suppose $\alpha>0$ (we have $\alpha<1$ by irreducibility), and that $A^{\prime}:=A/\alpha$ is irreducible and aperiodic. We claim that $k(\mu)=2$. This may be shown as follows.
Let $W=(W_{n}:n\geq 0)$ be a Markov chain on $S_{1}$ with transition matrix $A^{\prime}$, and let $i,j\in S_{1}$. By Theorem 2.5, if $W^{i}$ and $W^{j}$ are two independent versions of $W$, starting respectively at $i$ and $j$, then there exists $M<\infty$ such that
$\mathbb{P}(W_{m}^{i}=W_{m}^{j}\text{ for some }m\leq M)\geq\tfrac{1}{2}.$
It follows that
$\mathbb{P}(T_{i,j}<\infty)$ $\geq\mathbb{P}(T_{i,j}<\infty,\ \Pi_{m}=I\text{ for }m=1,2,\dots,M)$
$\geq\alpha^{M}\mathbb{P}(W_{m}^{i}=W_{m}^{j}\text{ for some }m\leq M)\geq\tfrac{1}{2}\alpha^{m}.$
[…]
Finally, in preparation for the next section we state a result from [8] which concerns the transition matrix
$P_{n}=\begin{pmatrix}n^{-1}&n^{-1}&\cdots&n^{-1}\\ n^{-1}&n^{-1}&\cdots&n^{-1}\\ \vdots&\vdots&\ddots&\vdots\\ n^{-1}&n^{-1}&\cdots&n^{-1}\end{pmatrix}$
on the state space $S=\{1,2,\dots,n\}$ with equal entries.
Theorem 5.5 ([8, Thm 7]).
For $n\geq 2$ there exists a block measure $\mu\in\mathcal{L}(P_{n})$ with $k(\mu)=\ell$ if and only if $\ell\mid n$. In particular, $K(P_{n})\supseteq\{\ell:\ell\mid n\}$. For $n\geq 3$, we have $n-1\notin K(P_{n})$.
6. Existence of non-block measures
Many of our measures $\mu\in\mathcal{L}_{P}$ so far have been block measures, though we encountered a certain $\mu$ in Example 3.10 which (by Theorem 4.4) is not a block measure. By Theorem 5.5, $n-1\notin K(P_{n})$, whence $K(P_{n})=\{\ell:\ell\mid n\}$ when $n\leq 4$. We do not know whether $K(P_{n})=\{\ell:\ell\mid n\}$ when $n>4$. In this section we shall construct a family of non-block measures for $P_{n}$.
Theorem 6.1.
Let $n\geq 4$ and suppose $\ell\neq 1$ and $\ell\mid n$. There exists a non-block measure $\mu\in\mathcal{L}_{P_{n}}$ with $k(\mu)=\ell$.
Before embarking on the proof of Theorem 6.1, we present an illustrative example.
Example 6.2.
Let $n=6$, $\ell=2$, and $b=n/\ell=3$. Let $S_{1}=\{1,2\}$, $S_{2}=\{3,4\}$, and $S_{3}=\{5,6\}$. Let $\widehat{\mathcal{F}}\subset\mathcal{F}_{S}$ consist of the six functions
$f_{1,1}$ $=(121212),$ $f_{1,2}$ $=(212121),$
$f_{2,1}$ $=(343443),$ $f_{2,2}$ $=(434334),$
$f_{3,1}$ $=(566565),$ $f_{3,2}$ $=(655656).$
We make the following observations.
  • (i)
    For each $u,v\in\{1,2,\dots,6\}$, $f(u)=v$ for exactly one $f\in\widehat{\mathcal{F}}$. Thus, taking $\mu$ to be uniform on $\widehat{\mathcal{F}}$, we obtain that $\mu\in\mathcal{L}_{P_{6}}$.
  • (ii)
    Let $f\in\widehat{\mathcal{F}}$. For each $i$ there exists $j$ such that $f$ maps $S_{i}$ onto $S_{j}$. Thus no pair of states in any $S_{i}$ will ever coalesce.
  • (iii)
[…]
Proof of Theorem 6.1.
Let $n,\ell$ be as in the statement of the theorem. We will imitate Example 6.2, and the reader may wish to refer back to that example. Let $b=n/\ell$ and
$S_{r}=\bigl\{(r-1)\ell+1,(r-1)\ell+2,\dots,r\ell\bigr\},\qquad r=1,2,\dots,b,$
noting that $|S_{r}|=\ell$ for each $r$. When considering $S_{r}$ as a vector rather than as a set, we order its elements in increasing order.
We will construct a set $\widehat{\mathcal{F}}$ comprising $n$ ($b\ell$) functions $f_{i,j}$ with $i\in\{1,2,\dots,b\}$, $j\in\{1,2,\dots,\ell\}$; $\mu$ will be the uniform measure on $\widehat{\mathcal{F}}$. Before giving a formal definition of $\widehat{\mathcal{F}}$, we summarise its main properties as follows:
  • (i)
    for every $u,v\in\{1,2,\dots,n\}$ there is exactly one $f\in\widehat{\mathcal{F}}$ such that $f(u)=v$,
  • (ii)
    for each $f\in\widehat{\mathcal{F}}$, there is a block $S_{f}$ such that every block $S_{r}$ is mapped by $f$ onto $S_{f}$ (viewed as sets),
  • (iii)
    the elements of each $S_{r}$ (viewed as a vector) are mapped to a certain permutation $\pi^{r}S_{f}$ of $S_{f}$.
We turn now to a formal definition, beginning with some notation. Following Definition 3.9, each $f_{i,j}$ can be expressed in the form $(x_{1}x_{2}\cdots x_{n})$, which is to say that $f_{i,j}(u)=x_{u}$. It is convenient to break such a vector into consecutive subsequences of length $\ell$, and we do this by adding vertical bars; thus, for example, we write $(x_{1}x_{2}\cdots x_{n})$ as
$\big(x_{1}\cdots x_{\ell}\big|x_{\ell+1}\cdots x_{2\ell}\big|\cdots\big|x_{(b-1)\ell+1}\cdots x_{b\ell}\big),$
and we term the $b$ subsequences therein ‘image blocks’. For each $f_{i,j}$, the image elements $x_{1},\dots,x_{\ell}$ are distinct; furthermore, when viewed as sets, we have that $\{x_{(i-1)\ell+1},\dots,x_{i\ell}\}=\{x_{1},\dots,x_{\ell}\}$ for each $i$, so that each subsequent image block of each $f_{i,j}$ is a permutation of the first image block of $f_{i,j}$.
We explain next the action of the $f_{i,j}$. Let $\rho$ denote the rotation permutation $i_{1}i_{2}\cdots i_{\ell}\mapsto i_{2}i_{3}\cdots i_{\ell}i_{1}$. We will express each $f_{i,j}$ in terms of a pair $(h_{i,j},{\bm{\pi}}^{i})$ where
  • (a)
    $h_{i,1}=S_{i}$ and $h_{i,j}=\rho^{j-1}S_{i}$ (with $S_{i}$ considered as a vector),
  • (b)
    for each $i=1,2,\dots,b$, there exists a vector ${\bm{\pi}}^{i}=(\pi^{i}_{2},\pi^{i}_{3},\dots,\pi^{i}_{b})$ of b1b-1 permutations (not necessarily distinct) of $\ell$ symbols,
  • (c)
    we set
    (6.1) $f_{i,j}=\big(h_{i,j}\big|\pi^{i}_{2}h_{i,j}\big|\cdots\big|\pi^{i}_{b}h_{i,j}\big).$
We discuss next how the ${\bm{\pi}}^{i}$ are defined, beginning with the simplest case $i=1$.
Let each $\pi^{1}_{j}$ be the identity permutation, so that, by (6.1),
$f_{1,1}=\big(h_{1,1}\big|\pi^{1}_{2}h_{1,1}\big|\cdots\big|\pi^{1}_{b}h_{1,1}\big)=\big(12\cdots\ell\big|12\cdots\ell\big|\cdots\big|12\cdots\ell\big).$
More generally, we let
(6.2) $f_{1,j}=\big(h_{1,j}\big|\pi^{1}_{2}h_{1,j}\big|\cdots\big|\pi^{1}_{b}h_{1,j}\big),$
where
$h_{1,j}=\pi_{m}^{1}h_{1,j}=\big(j(j+1)\cdots(\ell-1)\ell 12\cdots(j-1)\big),\qquad m\geq 2.$
We discuss next the case of $f_{i,j}$ with $i\geq 2$. Then $h_{i,1}=S_{i}$ and $h_{i,j}=\rho^{(j-1)}S_{i}$, and it remains to choose a suitable vector ${\bm{\pi}}^{i}$ of permutations. For $i=2,3,\dots,b$, let ${\bm{\pi}}^{i}=(\pi^{i}_{2},\dots,\pi^{i}_{b})$ be an ordered set of permutations satisfying $(\pi^{i}_{2},\dots,\pi^{i}_{b})\neq(\pi^{i^{\prime}}_{2},\dots,\pi^{i^{\prime}}_{b})$ for each $i^{\prime}<i$ (in other words, all the ordered sets of permutations will be distinct as ordered sets). We can always find distinct $\bm{\pi}^{1},\bm{\pi}^{2},\dots,\bm{\pi}^{b}$ since there are $(\ell!)^{b-1}$ distinct such vectors and $(\ell!)^{b-1}\geq 2^{b-1}\geq b$ (here we have used the fact that $b\geq 2$). The function $f_{i,j}$ is given by (6.1).
By construction, $\widehat{\mathcal{F}}$ has properties (i)–(iii) above. Let $\mu$ be the uniform probability measure on $\widehat{\mathcal{F}}$. By (i),
$\mu(\{f:f(u)=v\})=\frac{1}{n},\qquad u,v\in S,$
whence $\mu\in\mathcal{L}(P_{n})$.
We turn to the issue of coalescence. By (ii)–(iii) above, coalescence occurs at the first stage and not subsequently. Since there are at least two distinct $\bm{\pi}^{i}$, the coalescing classes are random (indeed the number of possibilities for the set of coalescing classes is the number of distinct $\bm{\pi}^{i}$, $i=1,2,\dots,b$). Since mappings between blocks are surjections, there are exactly $\ell$ coalescing classes $C_{1},C_{2},\dots,C_{\ell}$, and each such $C_{r}$ is a transversal of $\mathcal{S}$. Hence $k(\mu)=\ell$. ∎
Remark 6.3.
In the above proof, we have required that the ${\bm{\pi}}^{i}$ be distinct. It suffices for the proof that at least two of them are distinct. Suppose, on the contrary, that $\bm{\pi}^{j}=\bm{\pi}^{1}$ for all $j$. Then the set of coalescing classes is deterministic, and Theorem 4.4 applies. In particular since $\bm{\pi}^{1}$ is the identity permutation we may see that
$C_{j}=\{j,\ell+j,\dots,(b-1)\ell+j\},\quad j=1,2,\dots,\ell.$
[…]
7. Functions that generate transition matrices
In this final section we pose an inverse question. As usual, the state space is $S=\{1,2,\dots,n\}$. Recall the set $\mathcal{P}_{S}$ of transition matrices on $S$ and the set $\mathcal{F}_{S}$ of functions from $S$ to $S$. Given a subset $\mathcal{G}\subseteq\mathcal{F}_{S}$, let
$\mathcal{P}(\mathcal{G})=\{P\in\mathcal{P}_{S}:\exists\mu\in\mathcal{L}_{P}\text{ with }\text{\rm supp}(\mu)\subseteq\mathcal{G}\}.$
In other words, $\mathcal{P}(\mathcal{G})$ is the set of (irreducible) stochastic matrices that can be obtained by just using functions in $\mathcal{G}$.
Question 7.1.
For given $\mathcal{G}\subseteq\mathcal{F}_{S}$, what can be said about the set $\mathcal{P}(\mathcal{G})$?
Example 7.2.
For distinct $x,y\in S$ let
$f_{x,y}(v)=\begin{cases}y&\text{ if }v=x,\\ v&\text{ otherwise}.\end{cases}$
This is ‘almost’ the identity function (except that $x\mapsto y$). Let $\mathcal{G}=\{f_{x,y}:x,y\in S,\,x\neq y\}$, so that $\mathcal{G}$ contains n(n1)n(n-1) functions. Let $\bm{\alpha}=(\alpha_{x,y}:x\neq y)$ be a vector of non-negative reals satisfying
$\sum_{x,y:\,x\neq y}\alpha_{x,y}=1.$
Let $\mu_{\bm{\alpha}}$ be the probability measure on $\mathcal{G}$ that selects $f_{x,y}$ with probability $\alpha_{x,y}$. The corresponding $P=P_{\bm{\alpha}}=(p_{i,j})$ satisfies
$p_{i,j}=\mu\bigl(\{f:f(i)=j\}\bigr)=\mu(\{f_{i,j}\})=\alpha_{i,j},\qquad i\neq j,$
while
$p_{i,i}=\mu\bigl(\{f:f(i)=i\}\bigr)=1-\sum_{j:\,j\neq i}\alpha_{i,j}.$
[…]
Some very special cases admit precise answers to Question 7.1. Let $\mathcal{D}_{S}\subset\mathcal{P}_{S}$ denote the set of doubly stochastic matrices, and let $\mathcal{F}^{\text{\rm perm}}_{S}\subset\mathcal{F}_{S}$ denote the set of permutations of $S$.
Proposition 7.3.
  • (a)
    $\mathcal{P}(\mathcal{G})=\mathcal{D}_{S}$ if and only if $\mathcal{G}=\mathcal{F}^{\text{\rm perm}}_{S}$.
  • (b)
    $\mathcal{P}(\mathcal{G})=\mathcal{P}_{S}$ if and only if $\mathcal{G}=\mathcal{F}_{S}$.
Proof.
(a) This is a consequence of Theorem 2.6.
(b) Suppose first that $\mathcal{G}=\mathcal{F}_{S}$. For $P\in\mathcal{P}_{S}$ one can construct its independence coupling (recall Example 2.2). Therefore $\mathcal{P}(\mathcal{G})=\mathcal{P}_{S}$.
Suppose conversely that $\mathcal{G}$ is such that $\mathcal{P}(\mathcal{G})=\mathcal{P}_{S}$. We claim that for every $f\in\mathcal{F}_{S}$ there exists $P\in\mathcal{P}_{S}$ (hence $P\in\mathcal{P}(\mathcal{G})$) such that $f\in\text{\rm supp}(\mu)$ for every $\mu\in\mathcal{L}_{P}$. This implies that $f\in\mathcal{G}$ and therefore $\mathcal{G}=\mathcal{F}_{S}$ as required.
It remains to verify the above claim. Let $f=(j_{1}j_{2}\dots j_{n})\in\mathcal{F}_{S}$ (recall Definition 3.9). There exists $P_{f}=(p_{i,j})\in\mathcal{P}_{S}$ such that $p_{i,j_{i}}>1-1/n$ for every $i\in S$; the remaining terms $p_{i,j}$, $j\neq j_{i}$, are assumed strictly positive, thus implying that $P_{f}$ is irreducible. For $\mu\in\mathcal{L}_{P_{f}}$, we have
$\mu\bigl(\{g\in\mathcal{F}_{S}:g(i)=j_{i}\}\bigr)>1-\frac{1}{n},\qquad i=1,2,\dots,n,$
whence
$\mu(f)=\mu\left(\bigcap_{i\in S}\{g:g(i)=j_{i}\}\right)>0.$
Therefore, $f\in\text{\rm supp}(\mu)$. ∎
We recall the random transition matrix of Remark 2.4, with law $\mathbb{Q}$.
Question 7.4.
For given $\mathcal{G}\subset\mathcal{F}_{S}$, what can be said about $\mathbb{Q}(\mathcal{P}(\mathcal{G}))$?
This section closes with some partial answers to this question.
Firstly, in the proof of Proposition 7.3(b) it is shown that, if $f\notin\mathcal{G}$, then any matrix with $p_{i,f(i)}>1-n^{-1}$ for all $i\in S$ does not belong to $\mathcal{P}(\mathcal{G})$. It follows that $\mathbb{Q}(\mathcal{P}(\mathcal{G}))<1$ whenever $\mathcal{G}\neq\mathcal{F}_{S}$.
Secondly, in Example 7.2, we have $\mathbb{Q}(\mathcal{P}(\mathcal{G}))=0$ since $\mathcal{G}_{S}$ is the set of stochastic matrices with off-diagonal elements summing to $1$.
Thirdly, in the case $|S|=2$ there are just 4 functions, namely
$(11),(12),(21),(22).$
It is elementary that if $\mathcal{G}$ is missing two of them then $\mathbb{Q}(\mathcal{G})=0$. The following lemma shows that (when $n=2$) removing a single function one retains positive measure.
Proposition 7.5.
If $|S|=2$ and $f\in\mathcal{F}_{S}$, then $\mathbb{Q}(\mathcal{P}(\mathcal{F}_{S}\setminus\{f\}))=\frac{1}{2}$.
Proof.
Given a probability mass function $\bm{\alpha}$ on $\mathcal{F}_{S}$, the resulting matrix $P$ is given by
(7.1) $P=\begin{pmatrix}\alpha_{(11)}+\alpha_{(12)}&\alpha_{(21)}+\alpha_{(22)}\\ \alpha_{(11)}+\alpha_{(21)}&\alpha_{(12)}+\alpha_{(22)}\end{pmatrix}$.
Henceforth, fix $f=(uv)$ and write $u^{\prime}\neq u$ and $v^{\prime}\neq v$. Let $\mathcal{P}_{uv}$ denote the set of $P\in\mathcal{P}_{S}$ for which $p_{2,v}<p_{1,u^{\prime}}$. Since $p_{2,v}$ and $p_{1,u^{\prime}}$ are independent with the same probability density function (under $\mathbb{Q}$), we have that $\mathbb{Q}(\mathcal{P}_{uv})=\frac{1}{2}$. We claim that
(7.2) $\mathcal{P}(\mathcal{F}_{S}\setminus\{f\})=\mathcal{P}_{uv},$
and the proposition will follow immediately. It remains to prove (7.2).
Let $P\in\mathcal{P}(\mathcal{F}_{S}\setminus\{f\})$. There exists a mass function $\bm{\alpha}$} on $\mathcal{F}_{S}$ satisfying $\alpha_{(uv)}=0$ such that (7.1) holds; note that $\alpha_{(ab)}>0$ for $(ab)\neq(uv)$, $\mathbb{Q}$-a.s., and we will assume this. By (7.1),
$p_{2,v}=\alpha_{(u^{\prime}v)}<\alpha_{(u^{\prime}v)}+\alpha_{(u^{\prime}v^{\prime})}=p_{1,u^{\prime}}.$
Since $P$ is stochastic, we have also that $p_{2,v^{\prime}}>p_{1,u}$, and hence $P\in\mathcal{P}_{uv}$ as required.
Conversely, let $P\in\mathcal{P}_{uv}$, so that $p_{2,v}<p_{1,u^{\prime}}$. Set
$\alpha_{(uv)}=0,\quad\alpha_{(uv^{\prime})}=p_{1,u},\quad\alpha_{(u^{\prime}v)}=p_{2,v},\quad\alpha_{(u^{\prime}v^{\prime})}=1-(p_{1,u}+p_{2,v}),$
and note that
$1-\alpha_{(u^{\prime}v^{\prime})}=p_{1,u}+p_{2,v}<p_{1,u}+p_{1,u^{\prime}}=1.$
The stochastic matrix $Q=(q_{i,j})$ corresponding to $\bm{\alpha}$ (according to (7.1)) satisfies
$q_{1,u}$ $=\alpha_{(uv)}+\alpha_{(uv^{\prime})}=\alpha_{(uv^{\prime})}=p_{1,u},$
$q_{2,v}$ $=\alpha_{(uv)}+\alpha_{(u^{\prime}v)}=\alpha_{(u^{\prime}v)}=p_{2,v},$
whence $Q=P$. Therefore, $\mathcal{P}_{uv}\subseteq\mathcal{P}(\mathcal{F}_{S}\setminus\{f\})$ and the proof of (7.2) is complete. ∎
Finally, for any $\epsilon>0$, when $|S|$ is large there exist sets $\mathcal{G}\subseteq\mathcal{F}_{S}$ with size satisfying $|\mathcal{G}|>(1-\epsilon)|\mathcal{F}_{S}|$ for which $\mathcal{P}(\mathcal{G})$ has $\mathbb{Q}$-measure zero. Here are two examples.
Proposition 7.6.
We have $\mathbb{Q}(\mathcal{P}(\mathcal{G}))=0$ in the following two cases.
  • (a)
    Let $i,j\in S$ and $\mathcal{G}=\mathcal{G}_{i,j}=\{f\in\mathcal{F}_{S}:f(i)\neq j\}$.
  • (b)
    Let $i_{1},i_{2},j_{1},j_{2}\in S$ satisfy $i_{1}\neq i_{2}$, and let $\mathcal{G}$ be the set of all $f\in\mathcal{F}_{S}$ such that
    $f(i_{1})=j_{1}\quad\text{if and only if}\quad f(i_{2})=j_{2}.$
Let $n=|S|$. The cardinality of $\mathcal{G}$ in part (a) is $|\mathcal{F}_{S}|(1-n^{-1})$, and in part (b) is $|\mathcal{F}_{S}|(1-2n^{-1}+\text{\rm o}(n^{-1}))$.
Proof.
(a) For any $\mu$ supported on $\mathcal{G}_{i,j}$ we have $\mu(\{f:f(i)=j\})=0$, whence every $P=(p_{i,j})\in\mathcal{P}(\mathcal{G})$ has $p_{i,j}=0$. The probability of such $P$ satisfies $\mathbb{Q}(\mathcal{P}(\mathcal{G}))=0$.
(b) For any $\mu$ supported on such $\mathcal{G}$ we have
$\mu(\{f:f(i_{1})=j_{1}\})=\mu(\{f:f(i_{2})=j_{2}\}),$
so every $P=(p_{i,j})\in\mathcal{P}(\mathcal{G})$ has $p_{i_{1},j_{1}}=p_{i_{2},j_{2}}$. The claim follows. ∎
Example 7.7.
Let $n=3$, so that $|\mathcal{F}_{S}|=27$. Let $\mathcal{G}$ be the following set of 15 functions:
$\mathcal{G}$ $=\{(111),(311),(121),(321),\bm{(231)},$
$\qquad(112),(312),(122),(322),\bm{(232)},$
$\qquad(113),(313),(123),(323),\bm{(233)}\}.$
[…]
Acknowledgements
The authors are grateful to Tabet Aoul Alaa Eddine Islem for indicating an error in an earlier version of Theorem 4.3 as published in [8].